Sol Summary – January 10th

Crew 219 Sol Summary Report 10-01-2020

Sol: 5

Summary Title: Frozen Fog & Tang-glazed Cinnamon Rolls

Author’s name: Dave Masaitis

Mission Status: Nominal

Sol Activity Summary: The morning broke with the entire valley
shrouded in thick fog, dropping visibility next to nothing. By the
time EVA #9 was intended to depart, the morning sunshine had burned
off much of the fog, but ground conditions did not appear to support
safe maneuver conditions, so we decided to cancel it. The crew instead
busied themselves with science and maintenance, while H. Blackburn
prepared a wonderful lunch of pork sausage chili and fresh-baked
cornbread. The team for EVA #10 ate lunch and moved out on schedule,
achieving both their maneuver plan and sampling schedule. On the way
back, they recovered R. Raphael’s experimental mirror emplacement for
analysis. Upon their return they found another surprise, courtesy of
H. Blackburn – steaming hot chocolate and Tang-glazed cinnamon rolls –
a real treat! With full bellies, some crew members took a nap, while
others refined their reports for the evening Comms window. This
evening will consist of relaxed socialization while crew members
recoup for another Sol of hard work tomorrow.

Look Ahead Plan: We anticipate two EVA’s, and will continue ongoing
scientific analysis and maintenance operations.

Anomalies in work: None

Weather: Cold and foggy in the early morning, followed by clear,
sunny, and frozen conditions the rest of the day

Crew Physical Status: Nominal


EVA #9 – Cancelled at Commander’s Discretion

EVA #8 – Geologic Sample Collection South of Mid Ridge Planitia

Reports to be filed:

Sol Summary

EVA Report 10

EVA Requests 11 & 12

Operations Report

GreenHab Report

Research Report

Journalist’s Report

Photos (Including Photo of the Day)

Support Requested: Refer to Operations Report

Copyright © The Mars Society. All rights reserved. | Main Site